Grade 8Algebra

The sum of the ages of Anup and his father is 100.When Anup is as old as his father now,he will be 5 times as old as his son Anuj is now.Anuj will be 8 years older than Anup is now,when Anup is as old as his father.What are their ages now?

Sammir Malhotra
9 Years agoGrade 8

1 Answer

Vikas TU
9 Years ago
 Let period of Anup is a yrs 
along these lines, Father's age= (100-a) 
Anuj's age = (100-a)/5 
as indicated by given condition, 
(100-an)- a+(100-a)/5=a+8 
600-11a=5a+40 
16a=540 
a =35 yrs 
period of Anup is 35 yrs 
period of Anuj is 13yrs 
age of Anup's dad is 65yrs